# Diffusion Models: A Comprehensive Survey of Methods and Applications

### v14

@@ 2.3. Stochastic Differential Equations (Score SDEs)
| Xie et al. (Xie et al., 2022) | 2022 | Discrete | |||
| Gu et al. (Gu et al., 2022) | 2022 | Discrete | |||
@@ 5.1. Discrete Data
where is a one-hot column vector, and is the transition kernel of a lazy random walk. D3PM (Austin et al., 2021) accommodates discrete data in diffusion models by constructing the forward noising process with absorbing state kernels or discretized Gaussian kernels. Campbell et al. (2022) (Campbell et al., 2022) present the first continuous-time framework for discrete diffusion models. Leveraging Continuous Time Markov Chains, they are able to derive efficient samplers that outperform discrete counterparts, while providing a theoretical analysis on the error between the sample distribution and the true data distribution.
@@ 6.1. Large Language Models and Connections with Diffusion Models
Large Language Models (LLMs) (Brown et al., 2020; Anil et al., 2023; Achiam et al., 2023; Jiang et al., 2024; Yang et al., 2024e) have profoundly impacted the AI community, and showcased the advanced language comprehension and reasoning abilities. Recent works begin to extend their impressive reasoning abilities to visual generative tasks for overall generation planning. The collaboration between LLMs (ChatGPT, 2022; OpenAI, 2023b; Yang et al., 2024e) and diffusion models (Ramesh et al., 2022; Betker et al., 2023; Yang et al., 2023a, 2024h) can significantly improve the text-image alignment as well as the quality of generated images (Zhang et al., 2024a; Lian et al., 2023; Tian et al., 2024). For instance, RealCompo (Zhang et al., 2024a) utilizes LLMs to enhance the compositional generation of diffusion models by generating images grounded on bounding box layouts from the LLM. EditWorld (Yang et al., 2024f) composes a set of LLMs and pretrained diffusion models to generate an image editing dataset that contains numerous instructions with world knowledge (Ha and Schmidhuber, 2018). VideoTetris (Tian et al., 2024) uses the LLM to decompose text prompts along temporal axis for guiding video generation with smoother and more reasonable transitions. SemanticSDS (Yang et al., 2024g) and Trans4D (Zeng et al., 2024) extend the planning ability of LLMs to facilitate more complex 3D and 4D diffusion generation. Notably, RPG (Yang et al., 2024d) leverages the vision-language prior of multimodal LLMs to reason out complementary spatial layouts from text prompt, and manipulates the object compositions for diffusion models in both text-guided image generation and editing process, achieving SOTA performance in compositional synthesis scenarios and providing guidance for subsequent research.
@@ 7. Applications of Diffusion Models
| Multi-Modal Learning | Text-to-Image Generation | (Avrahami et al., 2022),(Ramesh et al., 2022),(Saharia et al., 2022b),(Nichol et al., 2022),(Gu et al., 2022),(Ruiz et al., 2022),(Kawar et al., 2022c),(Xu et al., 2022b),(Zhang and Agrawala, 2023),(Yang et al., 2023a),(Yang et al., 2024h),(Yang et al., 2024d),(Zhang et al., 2024b) |
@@ 7.3. Natural Language Generation
Natural language processing aims to understand, model, and manage human languages from different sources such as text or audio. Text generation has become one of the most critical and challenging tasks in natural language processing (Iqbal and Qureshi, 2020; Li et al., 2021a, b). It aims to compose plausible and readable text in the human language given input data (e.g., a sequence and keywords) or random noise. Numerous approaches based on diffusion models have been developed for text generation. Discrete Denoising Diffusion Probabilistic Models (D3PM) (Austin et al., 2021) introduces diffusion-like generative models for character-level text generation (Chelba et al., 2013). It generalize the multinomial diffusion model (Hoogeboom et al., 2021b) through going beyond corruption processes with uniform transition probabilities. Large autoregressive language models (LMs) is able to generate high-quality text (Brown et al., 2020; Chowdhery et al., 2022; Radford et al., 2019; Zhang et al., 2022b). To reliably deploy these LMs in real-world applications, the text generation process is usually expected to be controllable. It means we need to generate text that can satisfy desired requirements (e.g., topic, syntactic structure). Controlling the behavior of language models without re-training is a major and important problem in text generation (Keskar et al., 2019; Dathathri et al., 2019). Analog Bits (Chen et al., 2022) generates the analog bits to represent the discrete variables and further improves the sample quality with self-conditioning and asymmetric time intervals.
@@
Although recent methods have achieved significant successes on controlling simple sentence attributes (e.g., sentiment) (Krause et al., 2020; Yang and Klein, 2021), there is little progress on complex, fine-grained controls (e.g., syntactic structure). In order to tackle more complex controls, Diffusion-LM (Li et al., 2022a) proposes a new language model based on continuous diffusion. Diffusion-LM starts with a sequence of Gaussian noise vectors and incrementally denoises them into vectors corresponding to words. The gradual denoising steps help produce hierarchical continuous latent representations. This hierarchical and continuous latent variable can make it possible for simple, gradient-based methods to accomplish complex control. Similarly, DiffuSeq (Gong et al., 2023) also conducts diffusion process in latent space and proposes a new conditional diffusion model to accomplish more challenging text-to-text generation tasks. Ssd-LM (Han et al., 2022b) performs diffusion on the natural vocabulary space instead of a learned latent space, allowing the model to incorporate classifier guidance and modular control without any adaptation of off-the-shelf classifiers. CDCD (Dieleman et al., 2022) proposes to model categorical data (including texts) with diffusion models that are continuous both in time and input space, and designs a score interpolation technique for optimization.
@@ 7.4.3. Text-to-3D Generation
Modeling compositional 3D data distribution is a fundamental and critical task for generative models. Current feed-forward methods (Shue et al., 2023; Shi et al., 2024) are primarily capable of generating single objects and face challenges when creating more complex scenes containing multiple objects due to limited training data. Recently, a series of learnable-layout compositional methods have been proposed (Epstein et al., 2024; Vilesov et al., 2023; Han et al., 2024; Chen et al., 2024; Gao et al., 2024) . These methods combine multiple object-ad-hoc radiance fields and then optimize the positions of the radiance fields from external feedback. For example, Epstein et al. (2024) propose learning a distribution of reasonable layouts based solely on the knowledge from a large pre-trained text-to-image model. Vilesov et al. (2023) introduce an optimization method based on Monte-Carlo sampling and physical constraints. However, these forms of layout guidance are relatively coarse and not expressive enough for fine-grained control. Yang et al. (2024) (Yang et al., 2024g) address this problem by incorporating semantic embeddings that ensure view consistency and distinctly differentiate objects into SDS processes (namely SemanticSDS (Yang et al., 2024g)), which are flexible and expressive for optimizing 3D scenes. As shown in Fig. 15, SemanticSDS (Yang et al., 2024g) can achieve superior precision and quality in compositional text-to-3D generation over existing methods.
@@ 8. Future Directions
From Stable Diffusion to ChatGPT, Artificial Intelligence Generated Content (AIGC) has gained much attention in both academic and industrial circles. Generative Pre-Training is the core technique in GPT-1/2/3/4 (Radford et al., 2018, 2019; Ouyang et al., 2022; OpenAI, 2023a) and (Visual) ChatGPT (Wu et al., 2023), which exhibits promising generation performance and surprising emergent abilities (Wei et al., 2022) equipped with Large Language Models (LLMs) (Touvron et al., 2023) and Visual Foundation Models (Bommasani et al., 2021; Yu et al., 2022b; Yuan et al., 2021). It is interesting to transfer the generative pre-training (decoder-only) from GPT series to diffusion model class, evaluate the diffusion-based generation performance at scale, and analyse the emergent abilities of diffusion foundation models. Furthermore, combining LLMs with diffusion models have been proved to be a new promising direction (Yang et al., 2024d, f).

### v15

@@ 2.3. Stochastic Differential Equations (Score SDEs)
| Yang et al. (Yang et al., 2025a) | 2025 | Discrete | |||
| Wang et al. (Wang et al., 2025a) | 2025 | Discrete | |||
@@ 5.1. Discrete Data
where is a one-hot column vector, and is the transition kernel of a lazy random walk.
Building on discrete diffusion principles, MMaDA (Yang et al., 2025a) introduces a novel approach for multi-modal discrete data generation by incorporating masked autoencoder architectures within the diffusion framework. This method demonstrates improved performance on complex discrete sequences by leveraging cross-modal information. Similarly, TraceRL (Wang et al., 2025a) extends discrete diffusion models to more comprehensive reinforcement learning settings, where the discrete state-action sequences are generated through a specialized diffusion process that incorporates reward signals and policy constraints.
D3PM (Austin et al., 2021) accommodates discrete data in diffusion models by constructing the forward noising process with absorbing state kernels or discretized Gaussian kernels. Campbell et al. (2022) (Campbell et al., 2022) present the first continuous-time framework for discrete diffusion models. Leveraging Continuous Time Markov Chains, they are able to derive efficient samplers that outperform discrete counterparts, while providing a theoretical analysis on the error between the sample distribution and the true data distribution.
@@ 6.1. Large Language Models and Connections with Diffusion Models
Large Language Models (LLMs) (Brown et al., 2020; Anil et al., 2023; Achiam et al., 2023; Jiang et al., 2024; Yang et al., 2024e, 2025b; Zou et al., 2025; Wang et al., 2025b) have profoundly impacted the AI community, showcasing advanced language comprehension and reasoning abilities. Recent works have begun extending these impressive reasoning capabilities to visual generative tasks and overall generation planning, establishing powerful synergies between language understanding and multimodal generation.
The collaboration between LLMs (ChatGPT, 2022; OpenAI, 2023b; Yang et al., 2024e) and diffusion models (Ramesh et al., 2022; Betker et al., 2023; Yang et al., 2023a, 2024h) can significantly improve text-image alignment and the quality of generated content (Zhang et al., 2024a; Lian et al., 2023; Tian et al., 2024). For instance, RealCompo (Zhang et al., 2024a) utilizes LLMs to enhance the compositional generation of diffusion models by generating images grounded on bounding box layouts from the LLM. EditWorld (Yang et al., 2024f) composes a set of LLMs and pretrained diffusion models to generate an image editing dataset containing numerous instructions with world knowledge (Ha and Schmidhuber, 2018). VideoTetris (Tian et al., 2024) leverages LLMs to decompose text prompts along the temporal axis for guiding video generation with smoother and more reasonable transitions.
Beyond traditional text-to-image synthesis, recent advances have explored more sophisticated integrations. MMaDA (Yang et al., 2025a) demonstrates how LLMs can guide multi-modal discrete diffusion processes, where the language model’s understanding of cross-modal relationships enhances the generation of complex discrete sequences across different modalities. In reinforcement learning contexts, TraceRL (Wang et al., 2025a) combines the reasoning capabilities of LLMs with discrete diffusion models to generate optimal action sequences, where the language model provides high-level policy guidance while the diffusion process handles fine-grained action sampling.
Notably, RPG (Yang et al., 2024d) leverages the vision-language prior of multimodal LLMs to reason out complementary spatial layouts from text prompts, manipulating object compositions for diffusion models in both text-guided image generation and editing processes, achieving state-of-the-art performance in compositional synthesis scenarios.
These developments highlight a broader trend toward intelligent generation systems where LLMs serve as high-level planners and coordinators, while diffusion models handle the detailed generation process, creating more controllable and semantically coherent outputs across various domains and modalities.
@@ 7. Applications of Diffusion Models
| Multi-Modal Learning | Text-to-Image Generation | (Avrahami et al., 2022),(Ramesh et al., 2022),(Saharia et al., 2022b),(Nichol et al., 2022),(Gu et al., 2022),(Ruiz et al., 2022),(Kawar et al., 2022c),(Zhang and Agrawala, 2023),(Yang et al., 2023a),(Yang et al., 2024h),(Yang et al., 2024d),(Zhang et al., 2024b),(Yang et al., 2025a) |
@@ 7.3. Natural Language Generation
Natural language processing aims to understand, model, and manage human languages from different sources such as text or audio. Text generation has become one of the most critical and challenging tasks in natural language processing (Iqbal and Qureshi, 2020; Li et al., 2021a, b), aiming to compose plausible and readable text in human language given input data (e.g., a sequence and keywords) or random noise.
Numerous approaches based on diffusion models have been developed for text generation. Discrete Denoising Diffusion Probabilistic Models (D3PM) (Austin et al., 2021) introduces diffusion-like generative models for character-level text generation (Chelba et al., 2013), generalizing the multinomial diffusion model (Hoogeboom et al., 2021b) by going beyond corruption processes with uniform transition probabilities. Analog Bits (Chen et al., 2022) generates analog bits to represent discrete variables and further improves sample quality with self-conditioning and asymmetric time intervals.
While large autoregressive language models (LMs) can generate high-quality text (Brown et al., 2020; Chowdhery et al., 2022; Radford et al., 2019; Zhang et al., 2022b), reliable deployment in real-world applications typically requires controllable text generation to satisfy desired requirements (e.g., topic, syntactic structure). Controlling language model behavior without re-training remains a major challenge in text generation (Keskar et al., 2019; Dathathri et al., 2019). Although recent methods have achieved significant success on controlling simple sentence attributes (e.g., sentiment) (Krause et al., 2020; Yang and Klein, 2021), there has been limited progress on complex, fine-grained controls (e.g., syntactic structure).
@@
To tackle more complex controls, Diffusion-LM (Li et al., 2022a) proposes a new language model based on continuous diffusion that starts with Gaussian noise vectors and incrementally denoises them into word-corresponding vectors. The gradual denoising steps help produce hierarchical continuous latent representations, enabling simple gradient-based methods to accomplish complex control. Similarly, DiffuSeq (Gong et al., 2023) conducts diffusion processes in latent space and proposes a new conditional diffusion model for challenging text-to-text generation tasks.
Recent advances have explored more specialized applications of diffusion-based language generation. MMaDA (Yang et al., 2025a) extends diffusion-based text generation to multi-modal settings, where textual content is generated in conjunction with other modalities through masked autoencoder-enhanced diffusion processes (in Fig. 10). In reinforcement learning contexts, TraceRL (Wang et al., 2025a) applies diffusion models to generate natural language descriptions of optimal action sequences, bridging the gap between policy learning and natural language explanation.
These developments demonstrate the versatility of diffusion models in natural language generation, from basic text synthesis to complex multi-modal and interactive applications, opening new avenues for controllable and context-aware language generation.
@@ 7.4.1. Text-to-Image Generation
Recent advances have also explored building upon diffusion-based language model (DLLM) foundations for enhanced multi-modal generation capabilities. MMaDA (Yang et al., 2025a) leverages pre-trained DLLM backbones and extends them with masked autoencoder architectures to achieve superior multi-modal discrete data generation. By building on established DLLM foundations, MMaDA demonstrates how existing diffusion language models can be effectively adapted and enhanced for complex multi-modal synthesis tasks, opening new possibilities for foundation model reuse and extension in the diffusion domain.
@@ 7.4.3. Text-to-3D Generation
Modeling compositional 3D data distribution is a fundamental and critical task for generative models. Current feed-forward methods (Shue et al., 2023; Shi et al., 2024) are primarily capable of generating single objects and face challenges when creating more complex scenes containing multiple objects due to limited training data. Recently, a series of learnable-layout compositional methods have been proposed (Epstein et al., 2024; Vilesov et al., 2023; Han et al., 2024; Chen et al., 2024; Gao et al., 2024) . These methods combine multiple object-ad-hoc radiance fields and then optimize the positions of the radiance fields from external feedback. For example, Epstein et al. (2024) propose learning a distribution of reasonable layouts based solely on the knowledge from a large pre-trained text-to-image model. Vilesov et al. (2023) introduce an optimization method based on Monte-Carlo sampling and physical constraints.
@@ 8. Future Directions
From Stable Diffusion to ChatGPT, Artificial Intelligence Generated Content (AIGC) has gained much attention in both academic and industrial circles. Generative Pre-Training is the core technique in GPT-1/2/3/4 (Radford et al., 2018, 2019; Ouyang et al., 2022; OpenAI, 2023a) and (Visual) ChatGPT (Wu et al., 2023), which exhibits promising generation performance and surprising emergent abilities (Wei et al., 2022) equipped with Large Language Models (LLMs) (Touvron et al., 2023) and Visual Foundation Models (Bommasani et al., 2021; Yu et al., 2022b; Yuan et al., 2021).
It is interesting to transfer the generative pre-training (decoder-only) from GPT series to diffusion model class, evaluate the diffusion-based generation performance at scale, and analyse the emergent abilities of diffusion foundation models. Recent advances have demonstrated the potential of building specialized diffusion foundation models for different domains and applications. MMaDA (Yang et al., 2025a) exemplifies this trend by leveraging diffusion-based language model foundations to create a multi-modal generation framework that exhibits emergent cross-modal reasoning capabilities. Similarly, TraceRL (Wang et al., 2025a) explores how diffusion foundation models can be adapted for reinforcement learning scenarios, generating coherent action sequences while maintaining the scalable properties of foundation models.
Furthermore, combining LLMs with diffusion models has been proved to be a new promising direction (Yang et al., 2024d, f), opening up possibilities for more sophisticated AIGC systems (Liu et al., 2025) that leverage the complementary strengths of both paradigms. These developments suggest that the future of AIGC lies not only in scaling individual model types but also in creating hybrid architectures that can harness different generative principles for enhanced performance and broader applicability.